OpenStudy (anonymous):

A laser produces 20.0 mW of red light. In 1 hour, the laser emits 2.29 X 10^20 photons. What is the wavelength of the laser?

OpenStudy (aaronq):

find the amount of energy produced in joules, then divide it by the amount of photons. Use plancks equation to find the wavelength \(E_{photon}=\dfrac{h*c}{\lambda}\)

OpenStudy (anonymous):

I dont know what 20mW is ?

OpenStudy (aaronq):

mW = milliwatts

OpenStudy (anonymous):

If possible can you explain the process because i dont know how to find the # of Joules

OpenStudy (aaronq):

1 watt = 1 J/s 1 watt is 1 joule per second

OpenStudy (anonymous):

ohhh!!!!

OpenStudy (aaronq):

haha

OpenStudy (anonymous):

i divided .02 by 2.29*10^23 and got an answer. then divide that answer by h*c. my answer is wrong

OpenStudy (aaronq):

20 mW = 0.2 W = 0.2 J/s in 1 hour, (60 seconds*60 minutes) there are 3600 seconds so 0.2 J/s*3600 s = 720 J

OpenStudy (anonymous):

so its (2.29E20)(6.626E-23)(3E8)/720 = answer*10^9 to get it in nm

OpenStudy (aaronq):

damn, i noticed i screwed up on the conversion, it should be: 20 mW = 0.02 W = 0.02 J/s 0.02 J/s*3600 s = 72 J but yeah the rest looks good the answer would be in meters

OpenStudy (anonymous):

no worries thanks alot. this website is crazy

OpenStudy (aaronq):

no problem dude